OpenStudy (anonymous):

the lenght of the room is 2 times the width and the total area of the room is 8 square meters. what is the lenght of the living room?

OpenStudy (anonymous):

Area = (length) * (breadth) or, 8 = (2 b) * b or, 8 = 2b^2 or b^2 = 4 so, b = 2 Hence, length is 2*2 =4 meters

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome. I would suggest that you try such problems at first on your own. This would hone your math abilities.